A portfolio consists of Stocks A and B and has an expected return of 11.6 percent. Stock A has an expected return of 17.8 percent while Stock B is expected to return 8.4 percent. What is the portfolio weight of Stock A?

29.87%

61.98%

32.58%

34.04%

67.42%
Answer #1

Answer:

Option-(d): \(34.04 \%\)

I return \(=(\) Weight of Stock- \(A \times\) Return of Stock- \(A)+(\) Weight of \(S\) tock- \(B \times\) Return of \(S\) tock-B \()\)

be ' \(z\) '. Then, Stock-B's weight will be '1-z'. So, Let total weight of the portfolio be 1 and

\(11.6 \%=(2 \times 17.8 \%)+((1-z) \times 8.4 \%)\)

\(11.6 \%=(2 \times 17.8 \%)+8.4 \%-(2 \times 8.4 \%)\)

\(11.6 \%-8.4 \%=(2 \times 17.8 \%)-(2 \times 8.4 \%)\)

\(11.6 \%-8.4 \%=(2 \times 17.8 \%)-(z \times 8.4 \%)\)

\(3.2 \%=2 \times 9.4 \%\)

\(z=3.2 \% / 9.4 \%\)

\(z=34.04 \%\)

Therefore, Portfolio weight of Stock A is \(34.04 \%\).